For the rearrangement of ammonium cyanate to urea in aqueous solution at 50 °C, the rate of the reaction is determined by measuring the disappearance of NH4NCO.

NH4NCO---->(NH2)2CO

At the beginning of the reaction, the concentration of NH4NCO is 1.08 M. After 225 min the concentration has dropped to 0.294 M.

What is the rate of the reaction? ________ (mol NH4NCO/L) /min
Answer #1

NH4NCO---->(NH2)2CO

Rate of reaction = change in concentration /time taken

                            = 1.08-0.294/225-0

                             = 0.786/225

                             = 0.0035M/min

                             = 3.5*10^-3M/min >>>>answer